Problem 5 (extra credit problem) Assume that the Marriott has the following joint distribution with the market return. Under Bad, Good and Great states (that come with the probabilities 4, and , respectively) the market returns next year are (-5%, 15% and 35%) and the price of Marriott next year is expected to be, respectively: (S40m, $50m and S60m). Right now, Marriott has price of S44.00m and no debt. The risk free rate is 6.13%. You may find using a spreadsheet helpful in solving this problem. a) Use todays price and future Marriotts price to compute the return on the Marriotts equity in all three states. b) Use probabilities of every state to compute the expected return on the Marriott equity and the expected market return. c) Compute the variance of the market return and its covariance with Marriotts returns. You carn look up the formula for variance and covariance or just use the built-in functions in the spreadsheet d) Find the beta of Marriott equity. Is it higher than the markets beta? e) Assuming the long-term historical market risk premium of 8.4% find the appropriate discount rate for Marriott. Do not use the expected market return to calculate the market risk premium. only if their return is not lower than one implied by their risk (measured with C Assuming that the risk of the project is similar to the firms risk, what the value of the project. f) Assume that CAPM holds. Would you recommend buying Marriotts shares? We buy assets APM). ne of its project. g) Marriott expects $30m cash flows per year growing at 1% forever, from 0
Answer #1
a) Bad Good Great
A Current Price($ million) 44 44 44
B Expected Price next year 40 50 60
C=(B/A)-1 Return on Marriott's Equity -0.090909 0.136364 0.363636
Return on Marriott's Equity(%) -9.09% 13.64% 36.36%
b State Probability Equity Return (Equity return)*(Probability) Market Return (Market return)*(Probability)
Bad 0.25 -9.09% -0.02273 -5% -0.0125
good 0.5 13.64% 0.068182 15% 0.075
Great 0.25 36.36% 0.090909 35% 0.0875
SUM 0.136364 SUM 0.1500
Expected Return on Marriott's Equity 0.1363636
Expected Return on Marriott's Equity(%) 13.64%
Expected Return on Market 0.1500
Expected Return on Market(%) 15.00%
c Variance of Market Return: p A B=(A-15) C=B^2 E=C*P
Probability Market Return(%) Deviation from expected (%) Deviation Squared (Deviation Squared)*(Probability)
0.25         (5.00)      (20.00)      400.00      100.00
0.5         15.00                -                  -                  -  
0.25         35.00         20.00      400.00      100.00
SUM      200.00
Variance of Market Return        200.00
Equity Return(%) Market Return(%)
-9.09         (5.00)
13.64         15.00
36.36         35.00
Covariance with Marriott's Return 303 (Using COVAR function of excel)
d Beta of Marriott's Equity
Re=Expected Return of Equity(%) 13.64
Rm=Expected Return of Market(%) 15.00
Rf=Risk Free Rate(%) 6.13
Re=Rf+Beta*(Rm-Rf)
13.64=6.13+Beta*(15-6.13)
Beta=(13.64-6.13)/(15-6.13) 0.8466742
Beta of Marriott's Equity 0.8466742
Please upload balance as a separate question